OpenStudy (anonymous):

Of three positive integers, the second is twice the first, and the third is twice the second. One of these integers is 17 more than another. What is the sum of the three integers?

OpenStudy (anonymous):

Problem: x + 2x + 2*2x = y. Where 2x - x = 17 or 4x - 2x = 17 or 4x - x = 17. x = 17 or 2x = 17 or 3x = 17, only the first of these is an integer, so x = 17. And the answer is: 17 + 2*17 + 4*17 = 119